INFRARED LASER SPECTROSCOPY OF THE HCN-SODIUM COMPLEX EMBEDDED IN A HELIUM NANODROPLET

\underline{G. E. DOUBERLY}, T. M. FALCONER, and R. E. MILLER, Department of Chemistry, The University of North Carolina at Chapel Hill, Chapel Hill, NC 27599.

It is well established that sodium atoms reside on the surface of helium droplets rather than being completely solvated upon pick-up. However, subsequent pick-up of a polar molecule, such as HCN or HF, results in the solvation of the sodium atom upon complexation with the chromophore. The infrared spectrum of the linear HCN-Na complex reveals an anomolous reduction in the rotational constant in comparison to the ab initio RCCSD(T) value. This result is rationalized in terms of the HCN-Na-helium intermolecular potential energy surface.